Circular No. 5612

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CYGNI 1992 C. E. Woodward, University of Wyoming; D. Van Buren, California

Institute of Technology; P. Hofner, University of Wisconsin, Madison; and M. A. Greenhouse, National Air and Space Museum, Smithsonian Institution, write: “We have obtained near-infrared spectra of N Cyg 1992 (spectral range 3.590-3.790 μm, resolution 0.0033 μm) using the 2.1-m telescope (+ CRSP) at Kitt Peak National Observatory on Sept. 9.29 UT. Prominent in the spectra of the nova is the coronal line [Al VI] (observed wavelength 3.6480 ± 0.0014 μm) and the hydrogen recombination line Pfγ (observed wavelength 3.7384 ± 0.0015 μm). The discovery of the [Al VI] line at this epoch suggests that the nova has entered into the coronal line phase of its evolution ∼ 200 days after outburst. This evolutionary timescale is comparable to that seen in other novae that have exhibited coronal line emission such as V1500 Cyg and QU Vul. Preliminary analysis of the N Cyg 1992 spectra yields the following line flux estimates (× 10^{−19} W cm^{−2}): [Al VI] = 1.10 ± 0.13; Pfγ = 1.18 ± 0.25.”

R. Monella, Covo, Italy, telexes the following photometry (comparison star BD +52^{o}2734): Sept. 1.888 UT, V = 9.27, B–V = +0.18; 3.917, 9.29, +0.18; 5.925, 9.28, +0.16; 6.968, 9.30, +0.17.

GEMINGA P. R. Vishwanath, G. P. Satyanarayana, P. V. Ramanamurthy, and

P. N. Bhat, Tata Institute of Fundamental Research, Bombay, report: “We have reanalyzed the TeV γ-ray data taken on Geminga with the atmospheric Čerenkov Technique at Ootacamund during 1984 Dec.-1985 Feb. (cf. Bhat et al. 1987, A.Ap. 171, 84). A total of 28 hr of data at an energy threshold of 0.8 TeV was analysed for the presence of the 237-ms period seen by the Rosat, EGRET, and COS-B telescopes. The phase analysis done with the ephemeris of Hermsen et al. (IAUC 5541) revealed the presence of two peaks separated by 0.5 in phase. The probability that the two peaks are due to chance is 0.1 percent. The two peaks coincide in absolute phase with the position of the two peaks in COS-B data (Hermsen, private communication). The flux of γ-rays is calculated to be (2.1 ± 0.8) × 10^{−11} photon cm^{−2} s^{−1} at an energy threshold of 0.8 TeV.”

1992 September 14 Daniel W. E. Green